# An efficient implementation for solving the all pairs minimax path problem in an undirected dense graph

### v3

@@ Abstract.
We provide an efficient $O(n^{2})$ implementation for solving the all pairs minimax path problem or widest path problem in an undirected dense graph. It is a code implementation of the Algorithm 4 (MMJ distance by Calculation and Copy) in a previous paper. The distance matrix is also called the all points path distance (APPD). We conducted experiments to test the implementation and algorithm, compared it with several other algorithms for solving the APPD matrix. Result shows Algorithm 4 works good for solving the widest path or minimax path APPD matrix. It can drastically improve the efficiency for computing the APPD matrix. There are several theoretical outcomes which claim the APPD matrix can be solved accurately in $O(n^{2})$ However, they are impractical because there is no code implementation of these algorithms. It seems Algorithm 4 is the first algorithm that has an actual code implementation for solving the APPD matrix of minimax path or widest path problem in $O(n^{2})$, in an undirected dense graph.
@@
copyright: nonedoi: 00.000/000_0isbn: 000-$0$-00-000/00/00conference: 000; $0$; Beijing, Chinajournalyear: $0$price: 00.00
@@ 1. Introduction
where $G$ is the undirected dense graph. $\Theta_{(i,j,G)}$ is the set of all paths from node $i$ to node $j$. $p$ is a path from node $i$ to node $j$, $max\_weight(p)$ is the maximum weight in path $p$. $\Phi$ is the set of all maximum weights.   $min(\Phi)$ is the minimum of Set $\Phi$ (Liu, 2023).
The distance can also be called the longest-leg path distance (LLPD) (Little et al., 2020) or Min-Max-Jump distance (MMJ distance) (Liu, 2023). The all pairs minimax path distances calculate the distance between each pair of points in a dataset $X$ or graph $G$ It is also called all points path distance (APPD) (Little et al., 2020). It is a matrix of shape $n\times n$. A dataset $X$ can be straightforwardly converted to a complete graph.
@@ 3. Implementation of the algorithm
Note the three embedded for-loops make it look like an $O(n^{3})$ algorithm, but it is actually an $O(n^{2})$ algorithm. Because when the variable $i$ in $Line~{}21$ is small, both tree1 and tree2 are of size $O(n)$; but when the variable $i$ is large, both tree1 and tree2 are of size $O(1)$. The final net effect is that the three embedded for-loops only access each cell of the APPD matrix only once. Therefore, it is an $O(n^{2})$ algorithm.
@@ 6.2. Using parallel programming
If speed is the main concern of calculating the APPD matrix, we can use parallel programming to accelerate Algorithm 4. Firstly, we can use different processors for traversing the tree1 and tree2 in $Line~{}25~{}and~{}26$ of Figure 1(b). Secondly, we can copy the minimum spanning tree (MST) to many processors. For the $n$th processor, we just remove the $n$ largest edges, obtaining the $n$th tree1 and tree2, traversing them, then fill in the corresponding positions of the APPD matrix that are decided by the $n$th tree1 and tree2.

### v4

@@ Abstract.
We provide an efficient $O(n^{2})$ implementation for solving the all pairs minimax path problem or widest path problem in an undirected dense graph. It is a code implementation of the Algorithm 4 (MMJ distance by Calculation and Copy) in a previous paper. The distance matrix is also called the all points path distance (APPD). We conducted experiments to test the implementation and algorithm, compared it with several other algorithms for solving the APPD matrix. Result shows Algorithm 4 works good for solving the widest path or minimax path APPD matrix. It can drastically improve the efficiency for computing the APPD matrix. There are several theoretical outcomes which claim the APPD matrix can be solved accurately in $O(n^{2})$ . However, they are impractical because there is no code implementation of these algorithms. It seems Algorithm 4 is the first algorithm that has an actual code implementation for solving the APPD matrix of minimax path or widest path problem in $O(n^{2})$, in an undirected dense graph.
@@
copyright: nonedoi: 00.000/000_0isbn: 000-0000-00-000/00/00conference: 000; 0000; Beijing, Chinajournalyear: 0000price: 00.00
@@ 1. Introduction
where $G$ is the undirected dense graph. $\Theta_{(i,j,G)}$ is the set of all paths from node $i$ to node $j$. $p$ is a path from node $i$ to node $j$, max_weight(p)max\_weight(p) is the maximum weight in path $p$. $\Phi$ is the set of all maximum weights.   $min(\Phi)$ is the minimum of Set $\Phi$ (Liu, 2023).
The distance can also be called the longest-leg path distance (LLPD) (Little et al., 2020) or Min-Max-Jump distance (MMJ distance) (Liu, 2023). The all pairs minimax path distances calculate the distance between each pair of points in a dataset $X$ or graph $G$ . It is also called all points path distance (APPD) (Little et al., 2020). It is a matrix of shape $n\times n$. A dataset $X$ can be straightforwardly converted to a complete graph.
@@ 3. Implementation of the algorithm
Note the three embedded for-loops make it look like an $O(n^{3})$ algorithm, but it is actually an $O(n^{2})$ algorithm. Because when the variable $i$ in Line21Line~21 is small, both tree1 and tree2 are of size $O(n)$; but when the variable $i$ is large, both tree1 and tree2 are of size $O(1)$. The final net effect is that the three embedded for-loops only access each cell of the APPD matrix only once. Therefore, it is an $O(n^{2})$ algorithm.
@@
A further exploration shows we can also compute the APPD matrix by first sorting the edges in ascending order (from small to large), then adding the edges one-by-one to an empty MST tree.  [^2] See the “Sorting the edges of the MST in ascending order.ipynb” file in https://github.com/mike-liuliu/Algorithm_4
@@ 6.2. Using parallel programming
If speed is the main concern of calculating the APPD matrix, we can use parallel programming to accelerate Algorithm 4. Firstly, we can use different processors for traversing the tree1 and tree2 in Line25and26Line~25~and~26 of Figure 1(b). Secondly, we can copy the minimum spanning tree (MST) to many processors. For the $n$th processor, we just remove the $n$ largest edges, obtaining the $n$th tree1 and tree2, traversing them, then fill in the corresponding positions of the APPD matrix that are decided by the $n$th tree1 and tree2.
The parallel accelerated version of Algorithm 4 (MMJ distance by Calculation and Copy), is called Algorithm 13 (APPD accelerated by parallel computing). Here [^6] https://github.com/mike-liuliu/Algorithm_4/tree/main/Algorithm_4_in_cpp_parallel_computing is C++ code implementations of four variants of Algorithm 13.